OpenStudy (anonymous):

Multiply: (m + 7) (m - 7). A. m^2 - 49 B. m^2 - 14 C. m^2 - 14m + 49 D. m^2 - 49m + 14

OpenStudy (hoblos):

apply the formula (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

That is also known as FOIL.

OpenStudy (anonymous):

soo... B?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the two 7ms cancel each other out

OpenStudy (anonymous):

you do know what FOIL is right?

OpenStudy (hoblos):

@Chasitie 7^2 = 49

OpenStudy (anonymous):

Yeah but dont give her the answer

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

i know i know. ok FOIL mean First Outer Inner Last. if you just use the FOIL meathod you will be able to figure it out.

OpenStudy (anonymous):

no it is not c and i was talking to hoblos

OpenStudy (anonymous):

oh sorry jacobspa.

OpenStudy (anonymous):

its good

OpenStudy (anonymous):

ok chasitie, use FOIL.

OpenStudy (anonymous):

yes that is the easiest way

OpenStudy (anonymous):

i use it ALL the time. and to me... its is the only way.